OpenStudy (anonymous):

Find the LCD of (x2-9) and (x2+3x+2)

OpenStudy (anonymous):

lcd?

OpenStudy (anonymous):

lowest term?

OpenStudy (anonymous):

1st expression = x^2-9 since, it is in the form of a^2-b^2,we factorixe it by fromula (a+b) (a-b) then we get, (x)^2 -(3)^2 =(x-3) (x+3) Now, 2nd expression= x^2+3x+2 =x^2+ (2+1)x+2 = x^2+2x+x+2 taking x 1st two terms anmd 1 as common in last 2 terms,we get = x(x+2)+1(x+2) =(x+2) (x+1) therefore, LCD = (x-3) (x+3) (x+2) (x+1)

OpenStudy (anonymous):

sory,,dre's an error..actualy its''takin x common in 1st 2 terms" hope u got it!!!